8勺 Asig ment1-Encel в 1 u 〒 田, 쇼, 뉘 M rge&C.nter. $. % , un Coedtonal Femass Linked Cell Sort , nmtting Tshe 1 The balance shect of FIN 4100-601 Bank is listed bclow. Market yiclds arc in parcnthescs, and amounts arc in millions Asscts Cash Fed funds (2.05%) 3-month T-bills (3.25%) 8-year T-bonds (6.50%) 5-year munis (7.20%) 7-month C&I loans (4.8%) 2-year C&I loans (41 5%) Fixed-rate mortgages (5-10%) Liabilitics Demand dcposits Savings accounts (0.5%) $30 150 200 250 50 200 275 $150 50 0.02 0.22 MMDAs (3.5%) 7.55 (no minimum balance requirement) 4.25 3-month CDs (3.2%) 0.55 5-year CDs (5%) 1.65 Fed funds (2%) 1.25 260 175 350 225 0.2 4.85 0.02 10 6-month commercial paper (maturing in 5 months) Fixed-rate mortgages (5.30%) (maturing in 5 years) Fixed-rate mortgages (5.40%) (maturing in 20 years) 0.48 (4.05%) Subordinated debl: 450 275 355 300 0.55 4.45 7-year fixed rate (6.25%) Total liabilities 100 6.65 S1,610 15 16 17 18.25 Premiscs and cquipmert Total asscts Equity Total liabilitics and c S645 19 20 (1) What is the duration gap? 21 (2) What is the changc in cquity valuc forccastcd from the duration valucs for a predictcd incrcasc in interest rates of 0.5 pcrccnt? 23 24 Assignment instruction | Q1 Repriany Gap u2 Mdturty Gap I ug UET & Lpl | 04 Urstion & Conventy | QS ΔMVE

1) What is the duration gap?

2) What is the change in equity value forecasted from the duration values for a predicted increase in interest rates of 0.5 percent?
Answer #1

2) What is the change in equity value forecasted from the duration values for a predicted increase in interest rates of 0.5 percent?

ΔMVfedfunds= 0.0205 x 0.005/1.0205 x 150m = $15066.1

ΔMVT-bills= 0.22 x0.005/1.0325 x 200m = $213075

ΔMVT-bonds= 7.55 x 0.005/1.0650 x 250m = $6661502

ΔMVmunis= 4.25 x 0.005/1.0720 x 50m = $991138

ΔMVC&Iloans= 0.55 x0.005/1.0480 x 200m = $524809

ΔMVC&Iloans= 1.65 x 0.005/1.0415 x 275m = $2178348

ΔMVfixed-ratemortgages= 0.48 x0.005/1.0510 x 450m = $1027592

ΔMVfixed-ratemortgages= 4.45 x0.005/1.0530 x 275m = $5810779

ΔMVfixed-ratemortgages= 18.25 x0.005/1.0540 x 355m= $30734108

=>ΔMVA = $48156517

ΔMVsavings=1.25 x0.005/1.0050 x 50m = $310,945

ΔMVCDs= 0.20 x 0.005/1.0320 x 175m = $169,574

ΔMVCDs= 4.85 x 0.005/1.0500 x 350m = $8,083,333

ΔMVfedfunds= 0.02 x 0.005/1.0200 x 225m = $22,059

ΔMVcommericalpaper= 0.55 x 0.005/1.0405 x 300m = $792,888

ΔMVsubordinatedebt= 6.65 x 0.005/1.0625 x 100m = $3,129,412

=>ΔMVL = $12508211.

equity value forecasted from the duration values for increase of 0.5 percent in interest rates on assets and liabilities

ΔMVE = ΔMVA – ΔMVL = $48156517 – ($2508211) = $35648306